OpenStudy (anonymous):

given f(x)= 4x^(2), g(x)= -5x-3, and h(x)= √(x ) find [(f+g)oh](x)

OpenStudy (anonymous):

\[f(x)=4x^2,g(x)=-5x-3,(f+g)(x)=4x^2-5x-3\] \[(f+g)\circ h(x)=(f+g)(h(x))=(f+g)(\sqrt{x})=4(\sqrt{x})^2-5\sqrt{x}-3\]

OpenStudy (anonymous):

you can clean it up a little and write \[4x^2-5\sqrt{x}-3\]